OpenStudy (anonymous):

f(x)= 2x/x+3 and g(x)=5/x find: (f o g)(x) ?

OpenStudy (anonymous):

@marsss

OpenStudy (anonymous):

hey @Mertsj :D

OpenStudy (mertsj):

f composite g means to replace each x in the function f with the function g

OpenStudy (anonymous):

2(5/x) / (5/x)+3 right ?

OpenStudy (anonymous):

f o g means f(g(x)) if g(x) = 5/x and f(x) = 2x/(x+3) Then f o g = 2(5/x)/(5/x + 3)

OpenStudy (mertsj):



OpenStudy (anonymous):

ok cool :) but I'm stuck from there...

OpenStudy (mertsj):